Ncert Solutions Maths class 12th
Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th
- 2.5k
Questions
- 0
Discussions
- 0
Active Users
- 65
Followers
Have a question about Ncert Solutions Maths class 12th? Ask our experts
Keep it short & simple. Type complete word. Avoid abusive language.
Let the mirror image of the point (a, b, c), with respect to the plane 3x – 4y + 12z + 19 = 0 be (a−6, β, γ). If a + b + c = 5, then 7β-9γ is equal to.
Let the mirror image of the point (a, b, c), with respect to the plane 3x – 4y + 12z + 19 = 0 be (a−6, β, γ). If a + b + c = 5, then 7β-9γ is equal to.
(x-a)/3 = (y-b)/ (-4) = (z-c)/12 = -2 (3a-4b+12c+19)/ (3²+ (-4)²+12²)
(x-a)/3 = (y-b)/ (-4) = (z-c)/12 = (-6a+8b-24c-38)/169
(x, y, z) = (a–6, β, γ)
(a-b)-a)/3 = (β-b)/ (-4) = (γ-c)/12 = (-6a+8b-24c-38)/169
(β-b)/ (-4) = -2
=> β = 8+b
=> 3a – 4b + 12c = 150 . (i)
a + b + c = 5
=> 3a + 3b + 3c = 15 . (i
Let f : R → R be a function defined by f(x) = 2e²ˣ / (e²ˣ + e). Then f(1/100) + f(2/100) + f(3/100) + . + f(99/100) is equal to.
Let f : R → R be a function defined by f(x) = 2e²ˣ / (e²ˣ + e). Then f(1/100) + f(2/100) + f(3/100) + . + f(99/100) is equal to.
f (x) = 2e²? / (e²? +e? ) and f (1-x) = 2e²? / (e²? +e¹? )
∴ f (x) + f (1-x) = 1/2
i.e. f (x) + f (1-x) = 2
∴ f (1/100) + f (2/100) + . + f (99/100)
Σ? f (x/100) + f (1-x/100) + f (1/2)
= 49 x 2 + 1 = 99
Let X be a random variable having binomial distribution B(7, p). If P(X = 3) = 5P(X = 4), then the sum of the mean and the variance of X is:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
Let X be a random variable having binomial distribution B(7, p). If P(X = 3) = 5P(X = 4), then the sum of the mean and the variance of X is:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
Given P (X=3) = 5P (X=4) and n=7
=>? C? p³q? = 5? C? p? q³
=> q = 5p and also p + q = 1
=> p = 1/6 and q = 5/6
Mean = 7/6 and variance = 35/36
Mean + Variance
= 7/6 + 35/36 = 77/36
Five numbers x₁, x₂, x₃, x₄, x₅ are randomly selected from the numbers 1, 2, 3,., 18 and are arranged in the increasing order (x₁ < x₂ < x₃ < x₄ < x₅). The probability that x₂ = 7 and x₄ = 11 is:
Five numbers x₁, x₂, x₃, x₄, x₅ are randomly selected from the numbers 1, 2, 3,., 18 and are arranged in the increasing order (x₁ < x₂ < x₃ < x₄ < x₅). The probability that x₂ = 7 and x₄ = 11 is:
Total case = ¹? C?
Favourable cases
? C? (Select x? )
³C? (Select x? )
? C? (Select x? )
P = (6.3.7)/¹? C? = 1/68
If two straight lines whose direction cosines are given by the relations l+m-n=0, 3l2 +m2 +cnl=0 are parallel, then the positive value of c is:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
If two straight lines whose direction cosines are given by the relations l+m-n=0, 3l2 +m2 +cnl=0 are parallel, then the positive value of c is:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
l + m – n = 0 => n = l + m
3l² + m² + cnl = 0
3l² + m² + cl (l+m) = 0
= (3+c) (l/m)² + c (l/m) + 1 = 0
? Lines are parallel
D = 0
c = 4 (as c > 0)
Let a be an integer such that lim (x→a) [18 - [1-x]] / [x-3a] exists, where [t] is greatest integer ≤ t. Then a is equal to:
Let a be an integer such that lim (x→a) [18 - [1-x]] / [x-3a] exists, where [t] is greatest integer ≤ t. Then a is equal to:
lim? (x→7) (18- [1-x])/ ( [x-3a])
exist & a∈I.
= lim? (x→7) (17- [-x])/ ( [x]-3a)
exist
RHL = lim? (x→7? ) (17- [-x])/ ( [x]-3a) = 25/ (7-3a) [a ≠ 7/3]
LHL = lim? (x→7? ) (17- [-x])/ ( [x]-3a) = 24/ (6-3a) [a ≠ 2]
LHL = RHL
25/ (7-3a) = 8/ (2-a)
∴ a = -6
Let A(1,0), B(6,2) and C(3/2, 6) be the vertices of a triangle ABC. If P is a point inside the triangle ABC such that the triangles APC, APB and BPC have equal areas, then the length of the line segment PQ, where Q is the point (-7/6, -1/3) is
Let A(1,0), B(6,2) and C(3/2, 6) be the vertices of a triangle ABC. If P is a point inside the triangle ABC such that the triangles APC, APB and BPC have equal areas, then the length of the line segment PQ, where Q is the point (-7/6, -1/3) is
P will be the centroid of triangle ABC.
The centroid P is (x? +x? +x? )/3, (y? +y? +y? )/3).
The coordinates of P are given as (17/6, 8/3).
The coordinates of Q are not given, but a calculation is shown.
PQ = √ (24/6)² + (9/3)²) = √ (4² + 3²) = √ (16+9) = √25 = 5.
This implies the coordinates of Q are s
If the normal to the curve y(x) = ∫(from 2 to x) (2t² - 15t + 10)dt at a point (a, b) is parallel to the line x + 3y = -5, a > 1, then the value of |a + 6b| is equal to.
If the normal to the curve y(x) = ∫(from 2 to x) (2t² - 15t + 10)dt at a point (a, b) is parallel to the line x + 3y = -5, a > 1, then the value of |a + 6b| is equal to.
y (x) = ∫? (2t² - 15t + 10)dt
dy/dx = 2x² - 15x + 10.
For tangent at (a, b), slope is m = dx/dy = 1 / (dy/dx) = 1 / (2a² - 15a + 10).
Given slope is -1/3.
2a² - 15a + 10 = -3
2a² - 15a + 13 = 0 (The provided solution has 2a²-15a+7=0, suggesting a different problem or a typo)
Following the image: 2a² - 15a
A pack of cards has one card missing. Two cards are drawn randomly and are found to be spades. The probability that the missing card is not a spade, is :
A pack of cards has one card missing. Two cards are drawn randomly and are found to be spades. The probability that the missing card is not a spade, is :
Two drawn cards are spades. There are 50 cards left.
The missing card could be a spade or not a spade.
P (missing card is spade) = 11/50 (since 11 spades remain out of 50 cards).
P (missing card is not spade) = 1 - 11/50 = 39/50
Let P be a plane lx + my + nz = 0 containing the line, (1 - x)/1 = (y + 4)/2 = (z + 2)/3. If plane P divides the line segment AB joining points A(-3, -6, 1) and B(2, 4, -3) in ratio k : 1 then the value of k is equal to:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
Let P be a plane lx + my + nz = 0 containing the line, (1 - x)/1 = (y + 4)/2 = (z + 2)/3. If plane P divides the line segment AB joining points A(-3, -6, 1) and B(2, 4, -3) in ratio k : 1 then the value of k is equal to:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
